# MissionGNN: Hierarchical Multimodal GNN-based Weakly Supervised Video Anomaly Recognition with Mission-Specific Knowledge Graph Generation

### v2

@@ Abstract
In the context of escalating safety concerns across various domains, the tasks of Video Anomaly Detection (VAD) and Video Anomaly Recognition (VAR) have emerged as critically important for applications in intelligent surveillance, evidence investigation, violence alerting, etc. These tasks, aimed at identifying and classifying deviations from normal behavior in video data, face significant challenges due to the rarity of anomalies which leads to extremely imbalanced data and the impracticality of extensive frame-level data annotation for supervised learning. This paper introduces a novel hierarchical graph neural network (GNN) based model MissionGNN that addresses these challenges by leveraging a state-of-the-art large language model and a comprehensive knowledge graph for efficient weakly supervised learning in VAR. Our approach circumvents the limitations of previous methods by avoiding heavy gradient computations on large multimodal models and enabling fully frame-level training without fixed video segmentation. Utilizing automated, mission-specific knowledge graph generation, our model provides a practical and efficient solution for real-time video analysis without the constraints of previous segmentation-based or multimodal approaches. Experimental validation on benchmark datasets demonstrates our model’s performance in VAD and VAR, highlighting its potential to redefine the landscape of anomaly detection and recognition in video surveillance systems.
I Introduction
@@
II Related Works
II-A Weakly Supervised Video Anomaly Detection and Recognition
@@
II-B Graph Neural Networks and Knowledge Graphs
@@
II-C Large Multimodal Models and Prompt Optimization
@@
III Methodology
III-A Overall Pipeline
@@
Details are discussed in the following subsections: we first address our mission-specific KG generation framework in subsection III-B, then the proposed hierarchical multimodal GNN model and short-term temporal model are introduced in subsection III-C and subsection III-D respectively, and finally training losses are described in subsection III-E.
III-B Mission-Specific Knowledge Graph Generation Framework
@@ III-C Hierarchical Multimodal GNN
III-D Short-term Temporal Model
@@
III-E Weakly-supervised Frame-level Training
@@
Sparsity and Smoothing Losses. We adapted the same two regularization loss terms that are used in [28]. First is done through sparsity loss term exploiting that the anomaly cases are rare in entire time frames. Sparsity loss tries to minimize the number of frames that are predicted as anomalies by following:
@@
IV Experiments
TABLE I: Memory requirements for computational graph over each model in MiB.
@@ IV Experiments
IV-A Experimental Settings
IV-A1 Implementation Details.
@@ IV-A Experimental Settings
IV-B Datasets
@@
IV-C Evaluation on Training Memory Efficiency
@@
To demonstrate the advantage in training in terms of memory consumption of our proposed approach, we measured GPU memory consumption specifically the amount of memory for computing computational graphs which is required to conduct gradient backpropagation. Table I shows measured GPU memory consumption used for computational graphs for each of ImageBind and our GNN model by different batch sizes. We can observe that our model consumes 2.66×2.$6$ to 5.45×5.45$\times$5.45 × lower memory. Considering previous approaches used learnable prompts or other types of gradient computation of the LLVs are also using additional models to process the output from the LLV, the memory efficiency can be increased. It enables memory-efficient training while maintaining or outperforming previous models in some benchmarks that are shown in subsection IV-D.
IV-D Evaluation on VAD and VAR
@@
Table II and Table III show the evaluation results of both VAD and VAR tasks on UCF-Crime dataset and XD-Violence dataset respectively. MissionGNN surpasses the performance of cutting-edge VAR tasks in the UCF-Crime dataset, albeit showing a slight lag in VAD tasks. In contrast, AnomalyCLIP exhibits superior performance in VAR tasks within the XD-Violence dataset, while our framework notably outperforms state-of-the-art VAD tasks in the same dataset.
IV-E Evaluation on Real-time VAD and VAR
@@
A key advantage of the MissionGNN framework is its ability to operate with only 30 consecutive frames during both training and inference, making it well-suited for real-time anomaly detection and recognition. In contrast, AnomalyCLIP, which achieves higher performance in some evaluations, requires the entire video sequence for both training and inference, limiting its practical use in real-time applications such as intelligent surveillance. To illustrate this distinction, we conducted a comparative evaluation on the UCF-Crime and XD-Violence datasets, constraining AnomalyCLIP to the same 30-frame input as our model. The results, shown in Table II and Table III, highlight the impact of this limitation on AnomalyCLIP’s performance. In the VAR task, restricting the frame count caused a significant performance drop across all anomaly classes for AnomalyCLIP, resulting in reductions of 4.5% in mAUC and 14.29% in mAP scores across both datasets. This adjustment altered the performance comparison between the two methods. On the UCF-Crime dataset, MissionGNN exhibited a 6.23% improvement over AnomalyCLIP, further widening the performance gap. On the XD-Violence dataset, however, the gap narrowed by 3.71%, reducing the original 18% difference, which emphasizes the limitations of previous approaches in real-time settings. In the VAD task, MissionGNN outperformed AnomalyCLIP on both datasets, with gains of 1.13% in AUC and 32.91% in AP scores. These findings underscore the performance degradation of AnomalyCLIP when adapted for real-time analysis, as reflected in its lower AUC and AP scores on both datasets, reinforcing the practical advantages of our proposed framework for real-time anomaly detection and recognition.
TABLE II: Results of the state-of-the-art methods and MissionGNN on UCF-Crime
TABLE III: Results of the state-of-the-art methods and MissionGNN on XD-Violence
V Discussion

### v3

@@ Abstract
[…]
1 Introduction
@@
2 Background and Related Works
2.1 Weakly Supervised Video Anomaly Detection and Recognition
@@
2.2 Graph Neural Networks and Knowledge Graphs
@@
2.3 Large Multimodal Models and Prompt Optimization
@@
3 Methodology
3.1 Overall Pipeline
@@
Details are discussed in the following subsections: we first address our mission-specific KG generation framework in subsection 3.2, then the proposed hierarchical multimodal GNN model and short-term temporal model are introduced in subsection 3.3 and subsection 3.4 respectively, and finally training losses are described in subsection 3.5.
3.2 Mission-Specific Knowledge Graph Generation Framework
@@ 3.3 Hierarchical Multimodal GNN
3.4 Short-term Temporal Model
@@
3.5 Weakly-supervised Frame-level Training
@@
Sparsity and Smoothing Losses. We adapted the same two regularization loss terms that are used in [28]. First regularization is done through sparsity loss term exploiting that the anomaly cases are rare in the entire time frames. Sparsity loss tries to minimize the number of frames that are predicted as anomalies by following:
@@
4 Experiments
Table 1: Memory requirements for computational graph over each model in MiB.
@@ 4 Experiments
4.1 Experimental Settings
4.1.1 Implementation Details.
@@ 4.1 Experimental Settings
Table 2: Results of the state-of-the-art methods and MissionGNN on UCF-Crime
Table 3: Results of the state-of-the-art methods and MissionGNN on XD-Violence
4.2 Datasets
@@
4.3 Evaluation on Training Memory Efficiency
@@
To demonstrate the advantage in training in terms of memory consumption of our proposed approach, we measured GPU memory consumption specifically the amount of memory for computing computational graphs which is required to conduct gradient backpropagation. Table 1 shows measured GPU memory consumption used for computational graphs for each of ImageBind and our GNN model by different batch sizes. We can observe that our model consumes 2.66×2.$6$ to 5.45×5.45$\times$5.45 × lower memory. Considering previous approaches used learnable prompts or other types of gradient computation of the LLVs are also using additional models to process the output from the LLV, the memory efficiency can be increased. It enables memory-efficient training while maintaining or outperforming previous models in some benchmarks that are shown in subsection 4.4.
4.4 Evaluation on VAD and VAR
@@
Table 2 and Table 3 show the evaluation results of both VAD and VAR tasks on UCF-Crime dataset and XD-Violence dataset respectively. MissionGNN surpasses the performance of cutting-edge VAR tasks in the UCF-Crime dataset, albeit showing a slight lag in VAD tasks. In contrast, AnomalyCLIP exhibits superior performance in VAR tasks within the XD-Violence dataset, while our framework notably outperforms state-of-the-art VAD tasks in the same dataset.
4.5 Evaluation on Real-time VAD and VAR
@@
A key advantage of the MissionGNN framework is its ability to operate with only 30 consecutive frames during both training and inference, making it well-suited for real-time anomaly detection and recognition. In contrast, AnomalyCLIP, which achieves higher performance in some evaluations, requires the entire video sequence for both training and inference, limiting its practical use in real-time applications such as intelligent surveillance. To illustrate this distinction, we conducted a comparative evaluation on the UCF-Crime and XD-Violence datasets, constraining AnomalyCLIP to the same 30-frame input as our model. The results, shown in Table 2 and Table 3, highlight the impact of this limitation on AnomalyCLIP’s performance. In the VAR task, restricting the frame count caused a significant performance drop across all anomaly classes for AnomalyCLIP, resulting in reductions of 4.5% in mAUC and 14.29% in mAP scores across both datasets. This adjustment altered the performance comparison between the two methods. On the UCF-Crime dataset, MissionGNN exhibited a 6.23% improvement over AnomalyCLIP, further widening the performance gap. On the XD-Violence dataset, however, the gap narrowed by 3.71%, reducing the original 18% difference, which emphasizes the limitations of previous approaches in real-time settings. In the VAD task, MissionGNN outperformed AnomalyCLIP on both datasets, with gains of 1.13% in AUC and 32.91% in AP scores. These findings underscore the performance degradation of AnomalyCLIP when adapted for real-time analysis, as reflected in its lower AUC and AP scores on both datasets, reinforcing the practical advantages of our proposed framework for real-time anomaly detection and recognition.
5 Discussion
@@ Acknowledgements
Appendix
A. Ablation Study on KG Size Variations
Table 4 presents the results of our ablation study examining the impact of KG size variations on anomaly detection across different anomaly types. We explored two distinct scenarios: one where the KG is expanded by adding more key concept nodes, effectively widening the KG, and another where the KG is deepened through the incorporation of additional subgraphs from ConceptNet.
Our findings suggest that merely increasing the size of the KG can lead to overfitting, resulting in diminished performance. This underscores the importance of developing tailored, mission-specific KGs, as emphasized in our main paper, rather than simply enlarging existing KGs.
Table 4: Ablation study results on different size of KGs on XD-Violence dataset.
Case # of Key Concept Nodes Depth of Subgraphs AP - 30 1 98.42 Wide KG 60 1 97.66 Deep KG 30 2 97.06
B. Ablation Study on Decaying Threshold Method
Table 5 presents the results of our ablation study investigating the impact of various decay rates on the performance of our proposed decaying threshold-based anomaly localization method, specifically within the UCF-Crime dataset context. The findings demonstrate that our decaying threshold approach significantly enhances performance, evidenced by a 2.37% increase in the AUC score.
Moreover, the study highlights the critical importance of selecting an appropriate decay rate to maximize performance, with optimal results observed for decay rates $\alpha_{d}\geq 0.9$. This emphasizes the necessity of fine-tuning the decaying rate for achieving high performance in anomaly detection as we discuss in the main paper.
Table 5: Ablation study results on different decay rate of decaying threshold on UCF-Crime dataset.
Deacay Rate AUC 0.5 63.72 0.6 79.23 0.7 77.16 0.8 81.75 0.9 81.07 0.99 84.48 0.999 81.57 0.9999 83.57 1.0 (w/o Decaying Threshold) 82.11
C. Ablation Study on Loss Terms
Table 6 presents an ablation study evaluating the impact of each loss term in our proposed framework. The evaluation was conducted on the UCF-Crime dataset, measuring AUC scores after removing each loss term from the training loop. The results show that using all loss terms together yields the highest performance, with a significant score drop of at least 5.23% when any individual loss term is omitted. These findings highlight the critical importance of the proposed loss terms in achieving optimal model performance.
Table 6: Ablation study results on loss terms.
Case AUC Without $\mathcal{L}_{N}$ 60.49 Without $\mathcal{L}_{spa}$ 79.25 Without $\mathcal{L}_{s}mt$ 77.40 Full Model 84.48
Table 7: Ablation study results on different pre-trained joint embedding models.
Case Number of Supporting Modalities Joint Space Dimensionality AUC OpenCLIP (ViT-B-32) 2 512 84.12 ImageBind (imagebind_huge) 6 1024 84.48
Table 8: Ablation study results on a different number of frames.
Number of Frames AUC Inference GPU Memory Usage Inference Runtime 10 80.78 13,468 MiB 39 ms 20 83.47 13,478 MiB 43 ms 30 84.48 13,540 MiB 46 ms
Figure 3: Detailed process of mission-specific knowledge graph generation.
Figure 4: Portion of fails on generating valid (A) nodes and (B) edges on each layer by the number of attempts.
Figure 5: Example of KG for detecting the ”Shooting” category in the UCF-Crime dataset. Each color represents: Yellow: Sensor Node, Red: Key Concept Nodes, Blue: Sub-graph Nodes, Green: Encoding Node.
D. Ablation Study on Joint Embedding Models
Table 8 shows the results of our ablation study comparing different joint embedding models. We conducted the evaluation on the UCF-Crime dataset, measuring AUC scores using two joint embedding models: OpenCLIP and ImageBind. Aside from the embedding model and its dimensionality, all other hyperparameters were kept constant. The results demonstrate that our proposed framework performs consistently across both models, with only a 0.36% difference in scores. This suggests that the reasoning capability of our Hierarchical GNN framework is robust, regardless of the specific joint embedding model used.
E. Ablation Study on Local Context Size
Table 8 presents the results of our ablation study on different local context sizes. Using the UCF-Crime dataset, we measured AUC scores by varying the number of consecutive frames used during inference. The results show a trade-off between performance and resource usage, including GPU memory and runtime. As the local context size increases, the AUC scores improve, but at the cost of slightly higher memory and runtime demands. However, the performance gains tend to level off, with a smaller improvement between 20 and 30 frames compared to the larger gap between 10 and 20 frames. This suggests that after a certain context size, the model’s ability to benefit from additional frames diminishes, likely indicating diminishing returns on performance versus resource consumption.
F. Knowledge Graph Generation
Figure 3 illustrates the detailed pipeline for mission-specific knowledge graph (KG) generation. The process begins by taking the user’s input and initiating the Initial Reasoning Nodes Generation step. In this phase, initial keywords are generated using a large language model (LLM)—GPT-4 was employed in our evaluation—by utilizing pre-formatted prompts. Table 9 provides the prompts used in this step, where SUBJECT represents the user-defined mission, and DUP_NODES refers to keywords that have appeared in previous layers.
Following the generation of the initial reasoning nodes, the process moves to the Next Reasoning Nodes Generation step, where related keywords are inferred from the previously generated keywords. This step combines the LLM with ConceptNet to extract related keywords. ConceptNet’s output is used to guide the LLM in generating the next set of reasoning nodes. Table 10 presents the prompts used in this step, with COMMA_SEPARATED_LIST representing the previous layer’s keywords, and SUGGESTED_KEYWORDS containing related keywords extracted from ConceptNet.
After generating the next set of reasoning nodes, the system identifies logical relations between the nodes from the previous and current layers during the Reasoning Edge Generation step. Table 11 shows the prompts for this step, where SUBJECT is replaced with a newly generated keyword, COMMA_SEPARATED_LIST contains the keywords from the previous layer, and NOT_APPEARED_NODES refers to selected nodes that do not appear in the previous layer, as identified by the LLM.
Once the new layer and corresponding edges have been generated, the framework performs an Error Detection step to validate the correctness of each node and edge according to the definition of our reasoning KG. If any invalid nodes or edges are detected, the Error Correction step is triggered. During this step, errors are corrected using error correction prompts, with a maximum of three correction attempts allowed. If errors persist beyond the third attempt, the erroneous nodes or edges are pruned from the KG. After the error detection and correction loop, the framework returns to the Next Reasoning Nodes Generation step to expand the reasoning KG until the desired number of layers is achieved.
To evaluate the effectiveness of our error correction mechanism, we generated 190 reasoning KGs using our proposed method and tracked error rates by layer and by number of correction attempts, as illustrated in Figure 4. Across all layers, the number of errors significantly decreases with each correction attempt. For node generation (shown in Figure 4.(A)), the process results in a maximum of only 1% pruned nodes by the second layer. Similarly, edge generation (shown in Figure 4.(B)) results in a maximum of 0.02% pruned edges. These results demonstrate that the prompts used in our KG generation framework successfully facilitate knowledge extraction from both the LLM and ConceptNet. Figure 5 presents one KG we generated.
Table 9: Persona: Initial Reasoning Node Generation
Type of Prompt Prompt Format System Reference: A knowledge graph have hierarchical levels starting from naive observations to final prediction. Each level has the following inference words directly connected only with the previous level words. There are NOT the same words in different levels. Persona: You are a knowledge graph engineer who generates knowledge graph that will help to classify images. Objective: You will be provided a subject. Follow these steps to answer the user queries. Step 1. Observe 20 important words from a image which is related to the provided subject. Do not respond anything for this step. Step 2. Create a comma-separated list of the words that you observed. The comma-separated list you just created is first level of the knowledge graph. Keep in mind. Do not respond anything for this step. Step 3. Print first level of the knowledge graph on the first line. No extraneous text or characters other than the comma-separated list. User Subject: {SUBJECT} Error Correction The following concepts already appear in previous levels: {DUP_NODES} You must generate new concepts that can be inferred from previous level concepts. Correct this error and give a corrected answer. No extraneous text or characters other than the comma-separated list. Subject: {SUBJECT}
Table 10: Persona: Next Reasoning Nodes Generation
Type of Prompt Prompt Format System Reference: A knowledge graph have hierarchical levels starting from naive observations to final prediction. Each level has the following inference words directly connected only with the previous level words. There are NOT the same words in different levels. Persona: You are a knowledge graph engineer who generates knowledge graph that will help to classify images. Objective: You will be provided a subject. And you will be provided comma-separated list which is the previous level of the knowledge graph. And you will be provided suggested keywords. Follow these steps to answer the user queries. Step 1. Create words related to the provided subject which can be explained from combination of several words from previous level. Reference suggested keywords for this step. If you have better keywords, suggest them. Do not respond anything for this step. Step 2. Create a comma-separated list of the words that you just created in step 1. The length of comma-seperated list must be 20. The comma-separated list you just created is next level of the knowledge graph. Keep in mind. Do not respond anything for this step. Print next level of the knowledge graph on the first line. No extraneous text or characters other than the comma-separated list. User Subject: {SUBJECT} Comma-separated list: {COMMA_SEPARATED_LIST} Suggested keywords: {SUGGESTED_KEYWORDS} Error Correction The following concepts already appear in previous levels: {DUP_NODES} You must generate new concepts that can be inferred from previous level concepts. Correct this error and give a corrected answer. No extraneous text or characters other than the comma-separated list. Subject: {SUBJECT} Comma-separated list: {COMMA_SEPARATED_LIST} Suggested keywords: {SUGGESTED_KEYWORDS}
Table 11: Persona: Reasoning Edge Generation
Type of Prompt Prompt Format System Reference: A knowledge graph have hierarchical levels starting from naive observations to final prediction. Each level has the following inference words directly connected only with the previous level words. There are NOT the same words in different levels. Persona: You are a knowledge graph engineer who generates knowledge graph that will help to classify images. Objective: You will be provided a subject and a comma-separated list. Follow these steps to answer the user queries. Step 1. Select maximum 5 words from provided comma-separated list which are related to inferring provided subject. Do not respond anything for this step. Step 2. Create a comma-separated list of the selected words that you observed. Do not respond anything for this step. Step 3. Print the comma-separated list. No extraneous text or characters other than the comma-separated list. User Subject: {SUBJECT} Comma-separated list: {COMMA_SEPARATED_LIST} Error Correction The following concepts do not appear in the previous level nodes: {NOT_APPEARED_NODES} You must select concepts from the previous level concepts that can be important clues to infer the new concept. Correct this error and give a corrected answer. No extraneous text or characters other than the comma-separated list. Subject: {SUBJECT} Comma-separated list: {COMMA_SEPARATED_LIST}